OpenStudy (anonymous):

Solve x2 – 3x = –8 Answer

OpenStudy (mysesshou):

the square assumed \[x^2-3x=-8\] add 8 to each side of the = \[x^2-3x+8=0\] then..

OpenStudy (mysesshou):

factoring isn't simple, so you can use the quadratic equation.

OpenStudy (mysesshou):

do you know it?

OpenStudy (anonymous):

no but theres for muti choice questions want me to send them then maybe you can figure it out ? cuz i dony get at all

OpenStudy (anonymous):

*four * dont

OpenStudy (mysesshou):

\[x = \frac{-b \pm \sqrt(b^2-4*a*c)}{2*a}\]

OpenStudy (mysesshou):

\[x^2−3x+8=0\] \[x=\frac{-b \pm \sqrt{b^2-4*a*c}}{2*a} = x=\frac{-(-3) \pm \sqrt{(-3)^2-4*1*8}}{2*1}\]

OpenStudy (mysesshou):

\[x=\frac{3 \pm \sqrt{9-32}}{2}\]

OpenStudy (mysesshou):

\[x=\frac{3 \pm \sqrt{-23}}{2}\]

OpenStudy (mysesshou):

Then they pulled out the \[\sqrt{-1} = i\]

OpenStudy (mysesshou):

does that help any? confused anywhere?

OpenStudy (mysesshou):

so file called choice 2 is likely it.

OpenStudy (anonymous):

thank you

OpenStudy (mysesshou):

welcome !